Last class we reviewed GDP accounting-understanding GDP accounting helps us to better understand what GDP is measuring grow. Here is a simple example and helps us to understand how economics The farmer grows oranges and sells them to the retail store The retail store produces orange juice and sells if to consumers The retail store produces 200 gallons of juice (cost of $8/gallon) and sells 180 gallons to consumers for $10/gallon. The retail store pays the farmer $800 of which $700 is wage income and $100 is profit for the farmer The retail store pays wage income of S600 to its employees and rent of $200 on their store There are no other costs or expenses. Find the following values Consumption - what we sell to consumers so 180 x $10-$1,800 . Investment-inventory investment keeps track of goods produced but not sold and they are valued at cost because you do not know what price you will sell them for so 20x$8 $160 GDP C+IG+ NX so $1,960 Total expenditure = as above $1,960 Value added farmer S800 (the value of the oranges) Value added retail = value of output produced-purchased material so $1,960 $800 = $1,160 Total value added-S1,960 Wage income $1,300 . · Rental income= $200 . Profit income $100+$360 . Total income - $1,960 $460

I understand most of this EXCEPT, how the person got the INVESTMENT of 160. This is correct work, please show work on how this person got 160 for INVESTMENT.
Answer #1

Answer.

Total value of juice sold =180 x$ 10 =$1800.

To Farmer=$800.

So Consumption=1800-800=1000 dollars(Expenditure on intermediate products is excluded.The reason is that, that value is already included in the market price if the juice

Investment = wage+ rent +Inventory =600(wage)+200(rent)+160(20x8) =$960( we include the inventory in the investment)

GDP= C+ I +G+NX=1000+960 =$1960

Total expenditure=$1960(1000+960)

Value added by the farmer=$800

Value added by the retailer=1960-800=1160

Total value added=$1960

Wage income=$1300(700+600)

Rental income=$200

Profit income=100(farmer)+360(1160-800)=$460

Total income=$1960